OpenStudy (anonymous):

Simplify the following expression; express in terms of i. √-98

OpenStudy (angela210793):

=isqr98=7isqr2

OpenStudy (anonymous):

\[i \sqrt{2*49}=7i \sqrt{2}\]

OpenStudy (anonymous):

7i root2

OpenStudy (anonymous):

\[7i \sqrt{2}\]